OpenStudy (ladiesman217):

help please i dont understand this? Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a 6x^2 – 36x + 42 = 0

OpenStudy (radar):

Step 1: Divide thru by 6 getting: x^2 - 6x + 7 = 0 Step 2: add to both sides, the square of 1/2 of the x coefficient (6) 1/2 of 6 is 3, 3 squared is 9 the equation now will look like: x^2 - 6x + 9 = 9 Step 3. complying with the problem write the equation like they have requested (x - 3)^2 = 9

OpenStudy (radar):

Purpose of Step 1. Get the x^2 term with a coefficient of 1. purpose of step 2. Obtain a value that can be added that will result in the left side being a perfect square. Please note. I erred on this part. The equation would look like: X^2 - 6x + 7 + 9 = 9 Now subtract 7 from both sides getting: x^2 - 6x + 9 = 2 Now go with the last step (x - 3)^2 = 2

OpenStudy (radar):

Sorry about the error, it can happen.

OpenStudy (radar):

Do you have any further questions regarding this problem??

OpenStudy (ladiesman217):

no thanks